OpenStudy (gavin39):

Simplify completely quantity 4 x squared minus 7 x plus 3 all over x squared plus 5 x minus 6 quantity 4 x plus 1 over quantity x minus 1 quantity 4 x minus 1 over quantity x plus 6 quantity 4 x plus 3 over quantity x minus 1 quantity 4 x minus 3 over quantity x plus 6

OpenStudy (gavin39):

@hartnn @Mertsj

OpenStudy (mertsj):

\[\frac{4x^2-7x+3}{x^2+5x-6}=\frac{(x-1)(4x-3)}{(x+6)(x-1)}=\frac{4x-3}{x+6}\]